# Scaling Up Active Testing to Large Language Models

### v1

@@ A.3 Prompt formatting
A.4 Evaluation
For each input, the model outputs deterministic scores for each token in the vocabulary. We select the tokens corresponding to the labels of the dataset and apply the softmax function to obtain a final probability distribution over the possible labels. Any token that does not correspond to one of the labels is ignored.
A.5 Token selection
Tokenizers vary between models, and ideally, one token should correspond to one label, but this is not always the case. As highlighted by Kossen et al, [18], some models might represent task labels with multiple tokens. In such cases, we select only the relevant token that corresponds to the core of the word. For instance, the LLaMa-2 tokenizer encodes the word ‘objective’ as [12091] but ‘subjective’ as [4967, 573], which are tokens for ‘subject’ and ‘ive’. In this case, we select [12091] for ‘objective’ and [4967] for ‘subjective’.
A.6 In-context examples
@@
A.7 Sampling Details
To obtain deterministic, reproducible token generation, we set the maximum number of tokens to $1$, do not set a top-k nor top-p value and output the logits directly, from which we compute probabilities. As done in Kossen et al, [16], we clip values of the acquisition function to ensure that there is no zero-mass point and thus guarantee that the LURE is unbiased. In practice, all probabilities below $\alpha$ times the probability obtained by a uniform acquisition strategy are brought up to this limit value, where we set $\alpha=0.1$.
A.8 Computational resources
Our experiments are designed to be computationally efficient and therefore do not require substantial resources. Generating model predictions over the pool set is the main computational cost, typically requiring two GPUs for Llama-2 70B and one GPU for all other models used in this paper. Aside from this, the active testing procedure itself can be run efficiently on a small number of CPUs.
Appendix B Impact Statement
This paper presents work whose goal is to generally advance label- and compute-efficiency in model evaluation. As this is not limited to a particular area of application, it is difficult to evaluate the impact of this work. Generally speaking, we believe our work could play a role in reducing the computational costs, and thus the CO2 footprint, associated with model evaluation. As is often the case with research, we would caution against careless application of our method to real-world applications.
@@ Appendix C Additional results
C.1 Active evaluation of Llama models with Phi-2 and Gemma-3 1B surrogates
We further test active evaluation using even smaller surrogates: Phi-2 (2.7B, [13], MIT License) in Figure˜C.1 and Gemma-3 1B ([33], Gemma License) in Figure˜C.2. These models are limited in their capabilities. Although they generally perform well when evaluating 70b-zero or 7B-zero, they struggle with more performant few-shot configurations, leading to graceful failures comparable to random sampling. These results illustrate the limits of active evaluation with weak surrogates: effective active sampling depends on the surrogate providing informative signal, and when it does not, active testing offers little to no advantage over uniform sampling.
Figure C.1: Even very small surrogate models, here Phi-2, can achieve label-efficient evaluation in simple cases, although they struggle with evaluating our best performing target model, 70B-few.
Figure C.2: Even very small surrogate models, here Gemma-3 1B, can achieve label-efficient evaluation in simple cases, although they struggle with evaluating our best performing target model, 70B-few.
C.2 Weak surrogates perform worse on the more challenging MMLU dataset
[…]
Figure C.3: On the harder MMLU task, weak surrogates become insufficient. Yet, active evaluation with an uninformative surrogate fails gracefully, with performance close to uniform sampling and similar risk-estimation error decrease. This demonstrates the robustness of active testing in inefficient settings.
C.3 Examination of Pearson’s correlation coefficient
While we previously examined the failure of active evaluation on SST-2 with LURE-70B for the 70B-few configuration (Figure˜2), we also observe underperformance on Subj when using LURE-7B to evaluate 70B-few. Unlike the SST-2 case, filtering out high-loss points does not improve performance. To better understand this, we analyze in Table˜C.1 the Pearson correlation coefficient between the cross-entropy-based acquisition scores and the optimal acquisition function (negative-log likelihood of the target’s predictions). This correlation quantifies how well the surrogate approximates the optimal acquisition strategy. Notably, this is the only setting where the correlation is close to zero, and negative–indicating that the surrogate provides no meaningful guidance and the corresponding acquisition proposal may be selecting examples essentially at random. This explains why active evaluation performs similarly to uniform sampling in this case.
Table C.1: Pearson’s correlation coefficient between the cross-entropy of the surrogate’s and target’s predictions and the negative-log likelihood of the target’s predictions. Low or negative correlation indicates poor alignment between the surrogate and optimal acquisition function, explaining degraded active evaluation performance in specific cases. Active testing failure cases are shown in bold.
Model 70B-few 7B-few
Method LURE-Llama2-70B LURE-Llama2-7B LURE-Llama2-70B LURE-Llama2-7B
FPB $0.628$ $0.392$ $0.707$ $0.588$
SST-2 $\bf{0.285}$ $0.224$ $0.527$ $0.454$
Subj $0.364$ $\bf{-0.025}$ $0.802$ $0.652$
HS $0.461$ $0.146$ $0.555$ $0.384$
C.4 Risk-estimation values
In addition to figures shown in the main paper, we provide risk-estimation errors values for each experiment at acquisition steps $50,100,200,300$, and $400$ in Tables˜C.2, C.3, C.4, C.5 and C.6.
C.5 Evaluation results
Tables˜C.7 and C.8 provide the loss and accuracy numbers for Llama-2 7B, Llama-2 70B, and Gemma-3 4B, in both few- and zero-shot configurations, on our selection of target tasks. Results for Llama-2 here match those published in Kossen et al, [18].
Table C.2: Risk-estimation error for experiments on FPB using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $24.3373$ $11.5663$ $5.1944$ $3.4854$ $2.4122$
LURE-Llama2-7B $8.0511$ $3.6772$ $1.7589$ $1.0463$
LURE-Llama2-70B $\mathbf{4.2225}$ $\mathbf{2.0045}$ $\mathbf{0.98390}$ $\mathbf{0.64870}$
LURE-Gemma3-4B $7.1132$ $3.5176$ $1.6571$ $1.0540$
7B-zero Uniform random $9.4687$ $4.8302$ $2.1823$ $1.4524$ $1.0075$
LURE-Llama2-7B $2.5167$ $1.3094$ $0.59430$ $0.37380$
LURE-Llama2-70B $1.9294$ $0.96850$ $0.42460$ $0.27060$
LURE-Gemma3-4B $\mathbf{1.1421}$ $\mathbf{0.5849}$ $\mathbf{0.3681}$ $\mathbf{0.2341}$
70B-few Uniform random $10.5749$ $5.16310$ $2.48680$ $1.51150$ $0.977800$
LURE-Llama2-7B $6.3984$ $3.2058$ $1.3719$ $0.89110$ $0.63880$
LURE-Llama2-70B $5.0757$ $\mathbf{2.5057}$ $\mathbf{1.1921}$ $\mathbf{0.76740}$ $\mathbf{0.54300}$
LURE-Gemma3-4B $\mathbf{5.0390}$ $2.6267$ $1.3102$ $0.8344$ $0.5825$
7B-few Uniform random $14.3272$ $7.2276$ $3.4608$ $2.3147$ $1.5246$
LURE-Llama2-7B $8.9252$ $4.5853$ $2.3314$ $1.5486$ $1.1345$
LURE-Llama2-70B $\mathbf{4.8174}$ $\mathbf{2.1675}$ $\mathbf{1.1186}$ $\mathbf{0.65290}$ $\mathbf{0.45770}$
LURE-Gemma3-4B $6.5254$ $3.7684$ $1.9489$ $1.2371$ $0.8556$
Table C.3: Risk-estimation error for experiments on SST-2 using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $2.1154$ $1.0992$ $0.56060$ $0.34910$ $0.26150$
LURE-Llama2-7B $0.5386$ $0.3089$ $0.1551$ $0.1023$
LURE-Llama2-70B $\mathbf{0.4191}$ $\mathbf{0.2308}$ $\mathbf{0.1255}$ $\mathbf{0.08190}$
LURE-Gemma3-4B $0.5283$ $0.2790$ $0.1542$ $0.1042$
7B-zero Uniform random $12.4455$ $5.49$ $2.9545$ $1.9694$ $1.362$
LURE-Llama2-7B $2.7160$ $1.4229$ $0.66190$ $0.48410$
LURE-Llama2-70B $\mathbf{2.3740}$ $\mathbf{1.2437}$ $\mathbf{0.60510}$ $\mathbf{0.42530}$
LURE-Gemma3-4B $3.5371$ $1.6843$ $0.8699$ $0.6062$
70B-few Uniform random $36.6017$ $16.8825$ $9.4717$ $6.3474$ $4.3089$
LURE-Llama2-7B $\mathbf{24.8052}$ $\mathbf{14.7678}$ $\mathbf{7.9162}$ $\mathbf{5.2615}$ $\mathbf{4.0596}$
LURE-Llama2-70B $40.8407$ $30.077$ $21.3892$ $17.1808$ $14.7939$
LURE-Gemma3-4B $26.3377$ $16.8812$ $9.4499$ $7.1826$ $5.7175$
7B-few Uniform random $19.5177$ $9.5602$ $4.9347$ $3.1286$ $2.2497$
LURE-Llama2-7B $14.2747$ $8.3696$ $4.6265$ $3.2263$ $2.3764$
LURE-Llama2-70B $\mathbf{9.7216}$ $\mathbf{6.4914}$ $\mathbf{3.6886}$ $2.8971$ $2.1722$
LURE-Gemma3-4B $13.3504$ $7.0426$ $3.9896$ $\mathbf{2.8028}$ $\mathbf{1.9958}$
Table C.4: Risk-estimation error for experiments on Subjectivity using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $1.2104$ $0.64740$ $0.30890$ $0.21120$ $0.15430$
LURE-Llama2-7B $0.4688$ $0.2311$ $0.1113$ $0.07570$
LURE-Llama2-70B $\mathbf{0.139}$ $\mathbf{0.0758}$ $\mathbf{0.0373}$ $\mathbf{0.0251}$
LURE-Gemma3-4B $0.3809$ $0.1897$ $0.0887$ $0.0638$
7B-zero Uniform random $9.7212$ $5.1028$ $2.4378$ $1.5612$ $1.1429$
LURE-Llama2-7B $3.4150$ $1.6835$ $0.87130$ $0.58730$
LURE-Llama2-70B $\mathbf{1.6958}$ $\mathbf{0.89970}$ $\mathbf{0.44150}$ $\mathbf{0.29280}$
LURE-Gemma3-4B $2.8896$ $1.4536$ $0.7144$ $0.4787$
70B-few Uniform random $15.5609$ $6.6315$ $3.7465$ $2.5614$ $2.1347$
LURE-Llama2-7B $16.4075$ $8.2611$ $4.4942$ $3.1802$ $2.1987$
LURE-Llama2-70B $\mathbf{8.2473}$ $\mathbf{5.0366}$ $\mathbf{3.0052}$ $\mathbf{2.1589}$ $\mathbf{1.6293}$
LURE-Gemma3-4B $12.615$ $6.7789$ $3.3058$ $2.2544$ $1.7645$
7B-few Uniform random $7.4461$ $3.9287$ $1.8132$ $1.1308$ $0.87000$
LURE-Llama2-7B $5.2522$ $2.6268$ $1.2944$ $0.84500$ $0.63670$
LURE-Llama2-70B $\mathbf{1.8539}$ $\mathbf{1.0921}$ $\mathbf{0.62250}$ $\mathbf{0.42070}$ $\mathbf{0.30200}$
LURE-Gemma3-4B $4.1246$ $2.1545$ $1.0256$ $0.6868$ $0.4941$
Table C.5: Risk-estimation error for experiments on Hate Speech using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $2.3826$ $1.1493$ $6.3630$ $0.39690$ $0.30590$
LURE-Llama2-7B $0.8497$ $0.4371$ $0.2247$ $0.1524$
LURE-Llama2-70B $\mathbf{0.5703}$ $\mathbf{0.2996}$ $\mathbf{0.1418}$ $\mathbf{0.09900}$
LURE-Gemma3-4B $0.7508$ $0.3700$ $0.1822$ $0.1330$
7B-zero Uniform random $4.3455$ $\mathbf{2.3304}$ $\mathbf{1.3545}$ $0.96420$ $0.80470$
LURE-Llama2-7B $3.8653$ $1.9395$ $1.1326$ $0.86310$
LURE-Llama2-70B $3.0157$ 1.7693{1.7693} $\mathbf{0.94190}$ $\mathbf{0.73000}$
LURE-Gemma3-4B $3.5508$ $1.6981$ $0.9918$ $0.7766$
70B-few Uniform random $24.5497$ $11.8593$ $5.8947$ $4.263$ $2.897$
LURE-Llama2-7B $26.4253$ $13.5746$ $6.9755$ $4.249$ $3.0014$
LURE-Llama2-70B $19.8693$ $11.2704$ $6.1626$ $\mathbf{3.983}$ $2.9785$
LURE-Gemma3-4B $\mathbf{19.4658}$ $\mathbf{10.4853}$ $\mathbf{5.473}$ $\mathbf{3.6247}$ $\mathbf{2.767}$
7B-few Uniform random $23.7706$ $11.7404$ $6.6989$ $4.6215$ $3.4943$
LURE-Llama2-7B $19.6371$ $9.638$ $5.8037$ $4.0226$ $3.1452$
LURE-Llama2-70B $\mathbf{13.4897}$ $\mathbf{7.5196}$ $\mathbf{4.1388}$ $\mathbf{2.7654}$ $\mathbf{2.2691}$
LURE-Gemma3-4B $16.6135$ $8.7585$ $4.8131$ $3.4505$ $2.8044$
Table C.6: Risk-estimation error for experiments on MMLU using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $45.2305$ $22.0405$ $11.5532$ $7.2541$ $5.288$
LURE-Llama2-7B $57.137$ $28.4008$ $13.3008$ $8.4398$ $6.7655$
LURE-Llama2-70B $\mathbf{30.4079}$ 17.3273{17.3273} 9.5205{9.5205} 6.6253{6.6253} $4.9939$
LURE-Gemma3-4B $30.894$ $\mathbf{16.3988}$ $\mathbf{9.0229}$ $\mathbf{5.6838}$ $\mathbf{4.1935}$
7B-zero Uniform random $8.1557$ $4.1386$ $2.1924$ $1.4752$ $1.1799$
LURE-Llama2-7B $6.6835$ $3.2991$ $1.6763$ $1.2791$ $0.9767$
LURE-Llama2-70B $\mathbf{4.3678}$ $\mathbf{2.6164}$ $\mathbf{1.3063}$ $\mathbf{0.9088}$ $\mathbf{0.7330}$
LURE-Gemma3-4B $4.9507$ $2.6543$ $1.4700$ $0.9569$ $0.7622$
70B-few Uniform random 80.9096{80.9096} 41.8724{41.8724} 20.7615{20.7615} 12.5686{12.5686} 8.9944{8.9944}
LURE-Llama2-7B $94.483$ $51.3035$ $27.2732$ $17.5476$ $11.713$
LURE-Llama2-70B $84.9646$ $45.9222$ $23.3469$ $15.7314$ $12.5486$
LURE-Gemma3-4B $\mathbf{71.571}$ $\mathbf{34.8396}$ $\mathbf{19.3527}$ $\mathbf{12.9035}$ $\mathbf{9.2558}$
7B-few Uniform random $20.5838$ $10.142$ $4.8003$ $3.0952$ $2.4424$
LURE-Llama2-7B $16.2533$ $8.686$ $4.5997$ $3.047$ $2.2239$
LURE-Llama2-70B $\mathbf{7.242}$ $\mathbf{4.3427}$ $\mathbf{2.1991}$ $\mathbf{1.4639}$ $\mathbf{1.1166}$
LURE-Gemma3-4B $9.6443$ $4.9703$ $2.3827$ $1.694$ $1.2064$
Table C.7: Loss of the Llama-2 (L) 7B and 70B models and of the Gemma-3 (G) 4B model evaluated on the active set of various datasets.
L 7B-zero L 7B-few L 70B-zero L 70B-few G 4B-zero G 4B-few

### v2

@@ B.3 Prompt formatting
Then we introduce the sentence to be classified and ask for the corresponding label. The prompt is formatted as
@@
B.4 In-context examples
We randomly select in-context examples such that each class is represented in proportion to its actual proportion in the dataset, which we found improved the model’s accuracy. These examples are included in the input between the instruction and the prompt. Each input example is formatted as
[…]
All few-shot models are given 50 in-context examples for all datasets, except for the MMLU dataset, for which models receive only 5 examples due to token limit. Examples are ordered randomly once and fixed for all evaluations, ensuring that all models receive the same context. The set of in-context examples is therefore fixed beforehand and is common to both the target model and the surrogate model, for label efficiency and fair comparison.
B.5 Generating and processing model outputs
To obtain deterministic, reproducible token generation, we set the maximum number of tokens to 1, do not set a top-$k$ nor top-$p$ value and output the logits directly, from which we compute probabilities.
For each input, the model outputs a logit value for each token in the vocabulary. We select the logits corresponding to the labels of the dataset and apply the softmax function to obtain a final probability distribution over the possible labels. Any logit that does not correspond to one of the labels is ignored.
In cases where a model represents a task label with multiple tokens (kossen2024context), we select only the relevant token that corresponds to the core of the word. For instance, the Llama 2 tokeniser encodes the word “objective” as [12091] but “subjective” as [4967, 573], which are tokens for “subject” and “ive”. In this case we select [12091] for “objective” and [4967] for “subjective”.
B.6 Data acquisition
Following kossen2021active, we clip values of the acquisition function to ensure no zero-mass inputs and thus guarantee that the LURE is unbiased. All acquisition probabilities below $\alpha=0.1$ times the probability corresponding to uniform-random acquisition are brought up to this limit value.
B.7 Computational resources
Our experiments are designed to be computationally efficient and therefore do not require substantial resources. Generating model outputs over the pool set is the main computational cost, typically requiring two GPUs for Llama 2 70B and one GPU for all other models used in this work. Aside from this, the active-testing procedure itself can be run efficiently on a small number of CPUs.
@@ Appendix C Additional results
C.1 Model performance
Here we present accuracy and loss values for Llama2-7B, Llama2-70B and Gemma3-4B on the five datasets we study. Results for Llama 2 here match those published in kossen2024context.
Dataset L-7B-zero L-7B-few L-70B-zero L-70B-few G-4B-zero G-4B-few
FPB 66.75 90.60 26.25 94.15 29.70 91.00
SST-2 63.96 92.19 76.80 93.60 57.79 91.68
Subj 49.38 89.72 54.85 96.12 50.40 93.73
HS 11.38 89.28 84.18 90.23 42.87 89.42
MMLU 35.41 42.16 60.80 65.34 52.61 57.83
Table 2: Accuracy (%) of Llama 2 (L) 7B and 70B models and Gemma3-4B (G) model evaluated on pool sets.
Dataset L-7B-zero L-7B-few L-70B-zero L-70B-few G-4B-zero G-4B-few